# Intrinsic versus extrinsic cellular decision making

### v1

@@ II Data points
  • The cell seeding densities used in Polacheck et al. (2011) are $\rho=50$ mm-3 and $\rho=250$ mm-3. The cell seeding density used in Polacheck et al. (2014) is $\rho=600$ mm-3.
@@
  • The flow speeds applied in Polacheck et al. (2011) are $v=0.3$ $\mu$m/s and $v=3$ $\mu$m/s. The flow speed applied in Polacheck et al. (2014) is $v=4.6$ $\mu$m/s.
@@
  • At $\rho=50$ mm-3, cells migrated downstream on average for both flow speeds (see gray squares in Fig. 3B of Polacheck et al. (2011)). At $\rho=250$ mm-3, cells migrated upstream on average for both flow speeds (see gray circles in Fig. 3B of Polacheck et al. (2011)). At $\rho=600$ mm-3 and $v=4.6$ $\mu$m/s, cells migrated upstream on average (see lower left plot in Fig. 5C of Polacheck et al. (2014)).
For Fig. 3(b):
  • The flow speed applied in ran Moon et al. (2023) is $v=1.5$ $\mu$m/s.
@@
  • The gradient $g$ and average background concentration $c$ are computed between all pairs of data points in ran Moon et al. (2023) with (i) no flow (Fig. 2D in ran Moon et al. (2023)), (ii) a large downstream gradient (right region of Fig. 2F in ran Moon et al. (2023)), and (iii) a small upstream gradient (right region of Fig. 2E in ran Moon et al. (2023)). The values of $G=g\sqrt{a^{5}/c}$ are calculated with $a=10$ $\mu$m and averaged. The standard deviations in $G$ are smaller than the data points in Fig. 3(b).
@@
  • For the case of (iv) a small downstream gradient (left region of Fig. 2F in ran Moon et al. (2023)), because the concentration values are so small, we use the same procedure as above, but for the raw concentration data rather than the trial- and window-averaged data in Fig. 2F of ran Moon et al. (2023), for better accuracy. The error bar on this point in Fig. 3(b) is the standard deviation across six experimental trials of the average $G$ value in this region.
@@
  • For the above cases, cells migrated, on average, (i) up-gradient (Fig. 3C in ran Moon et al. (2023), second box plot), (ii) up-gradient (Fig. 3C in ran Moon et al. (2023), sixth box plot), (iii) nowhere (Fig. 3C in ran Moon et al. (2023), fifth box plot; compare with control), and (iv) upstream (Fig. 3C in ran Moon et al. (2023), seventh box plot). In the case of flow with no gradient, cells migrated upstream on average (Fig. 3C in ran Moon et al. (2023), third box plot).
@@
  • Typical cell surface concentrations of ephrin are on the order of $c_{2}=10^{2}$--$10^{3}$ $\mu$m-2, and ephrin receptors have been measured to respond to these concentrations (see Fig. 3 in Xu et al. (2011)). We use this range [error bars in Fig. 3(c)].
  • The diffusion constant of ephrin was measured at $D_{2}=2$ $\mu$m2/s (Fig. 2A in Xu et al. (2011)).
@@
  • For head-to-head cell-cell collisions, the EGF concentrations used in Lin et al. (2015) increase by $\Delta c=3.3$ nM on top of backgrounds of $c=4.1$ nM, $c=8.3$ nM, and $c=12.4$ nM (see Fig. 3c-e in Lin et al. (2015)). Only $f=40\%$ of this increase occurs within the cell microchannel, over a distance of approximately $d=250$ $\mu$m, and the profile is linear within the microchannel (see Fig. S1e of Lin et al. (2015)). Thus, $g=f\Delta c/d$. These three values of $G=g\sqrt{a^{5}/c}$ with $a=10$ $\mu$m are used in Fig. 3(c).
@@
  • The majority of cells migrated up-gradient (outcome 1 in Fig. 3a-e of Lin et al. (2015)) when $c=4.1$ nM and $c=8.3$ nM (see right bars in Fig. 3c and d of Lin et al. (2015)), whereas the majority of cells migrated away from the other cell (outcomes 2 and 3 in Fig. 3a-e of Lin et al. (2015)) when $c=12.4$ nM (see right bars in Fig. 3e of Lin et al. (2015)).

### v2

@@ Appendix C Data points
  • The cell seeding densities used in [19] are $\rho=50$ mm-3 and $\rho=250$ mm-3. The cell seeding density used in [20] is $\rho=600$ mm-3. The cell seeding density used in [29] is $\rho=250$ mm-3. The cell seeding density used in [23] is $\rho=300$ mm-3.
@@
  • The flow speeds applied in [19] are $v=0.3$ $\mu$m/s and $v=3$ $\mu$m/s. The flow speed applied in [20] is $v=4.6$ $\mu$m/s. The flow speed applied in [29] is $v=7$--$16$ $\mu$m/s (see Fig. 2D therein). The flow speed applied in [23] is $v=0.7$ $\mu$m/s or, in their pillar flow chamber, $v=0.2$--$0.8$ $\mu$m/s.
@@
  • In [19], at $\rho=50$ mm-3, cells migrated downstream on average for both flow speeds (see gray squares in Fig. 3B therein). In [19], at $\rho=250$ mm-3, cells migrated upstream on average for both flow speeds (see gray circles in Fig. 3B therein). In [20], cells migrated upstream on average (see lower left plot in Fig. 5C therein). In [29], distinct subpopulations of cells migrated downstream and upstream (see Fig. 5E therein), such that the total population did neither on average (see Fig. 3D therein, rightmost plot). In [23], cells polarized downstream in their radial flow chamber (see Fig. 4B therein) and migrated downstream in their pillar flow chamber (see Fig. 5F-H therein).
For Fig. 3(b), circles:
  • The flow speed applied in [15] is $v=1.5$ $\mu$m/s.
@@
  • The gradient $g$ and average background concentration $c$ are computed between all pairs of data points in [15] with (i) no flow (Fig. 2D in [15]), (ii) a large downstream gradient (right region of Fig. 2F in [15]), and (iii) a small upstream gradient (right region of Fig. 2E in [15]). The values of $G=g\sqrt{a^{5}/c}$ are calculated with $a=10$ $\mu$m and averaged. The standard deviations in $G$ are smaller than the data points in Fig. 3(b).
@@
  • For the case of (iv) a small downstream gradient (left region of Fig. 2F in [15]), because the concentration values are so small, we use the same procedure as above, but for the raw concentration data rather than the trial- and window-averaged data in Fig. 2F of [15], for better accuracy. The error bar on this point in Fig. 3(b) is the standard deviation across six experimental trials of the average $G$ value in this region.
@@
  • For the above cases, cells migrated, on average, (i) up-gradient (Fig. 3C in [15], second box plot), (ii) up-gradient (Fig. 3C in [15], sixth box plot), (iii) nowhere (Fig. 3C in [15], fifth box plot; compare with control), and (iv) upstream (Fig. 3C in [15], seventh box plot). In the case of flow with no gradient, cells migrated upstream on average (Fig. 3C in [15], third box plot).
For Fig. 3(b), triangles (details in Appendix D):
  • The flow speed applied in our experiments is $v=2.5$ $\mu$m/s.
  • In our experiments, cells are tracked across the entire chamber with width $L=1$ mm, excluding the microfluidic pillars that span $d=100$ $\mu$m on either side. Therefore, the average and standard deviation of $G$ across the chamber are computed as
[…]
    where $G(x)=g(x)\sqrt{a^{5}/c(x)}$, with [15]
    $c(x)$ $=(c_{2}-c_{1})\left(\frac{e^{vx/D}-1}{e^{vL/D}-1}\right)+c_{1},$ (15)
    $g(x)$ $=\frac{dc}{dx}=\frac{v(c_{2}-c_{1})e^{vx/D}}{D(e^{vL/D}-1)},$ (16)
    and $a=10$ $\mu$m. Here, $c_{1}$ and $c_{2}$ are the TGF-$\beta$ concentrations on the upstream and downstream sides of the chamber, respectively, and $D=150$ $\mu$m2/s is the TGF-$\beta$ diffusion coefficient.
  • Without flow ($v=0$), we used $c_{1}=0$ and $c_{2}=5$ nM, and cells migrated up-gradient on average [Fig. 4(g), red]. With flow: (i) when there was no gradient ($c_{1}=c_{2}=0$), cells migrated upstream on average [Fig. 4(g), blue]; (ii) when the gradient was small ($c_{1}=2.5$ nM, $c_{2}=5$ nM), cells migrated upstream on average [Fig. 4(g), purple]; and (iii) when the gradient was large ($c_{1}=0$, $c_{2}=10$ nM), cells migrated up-gradient on average [Fig. 4(g), brown].
@@
  • Typical cell surface concentrations of ephrin are on the order of $c_{2}=10^{2}$--$10^{3}$ $\mu$m-2, and ephrin receptors have been measured to respond to these concentrations (see Fig. 3 in [33]). We use this range for the error bars in Fig. 3(c).
  • The diffusion constant of ephrin was measured at $D_{2}=2$ $\mu$m2/s (Fig. 2A in [33]).
@@
  • For head-to-head cell-cell collisions, the EGF concentrations used in [22] increase by $\Delta c=3.3$ nM on top of backgrounds of $c=4.1$ nM, $c=8.3$ nM, and $c=12.4$ nM (see Fig. 3c-e in [22]). Only $f=40\%$ of this increase occurs within the cell microchannel, over a distance of approximately $d=250$ $\mu$m, and the profile is linear within the microchannel (see Fig. S1e of [22]). Thus, $g=f\Delta c/d$. These three values of $G=g\sqrt{a^{5}/c}$ with $a=10$ $\mu$m are used in Fig. 3(c).
@@
  • The majority of cells migrated up-gradient (outcome 1 in Fig. 3a-e of [22]) when $c=4.1$ nM and $c=8.3$ nM (see right bars in Fig. 3c and d of [22]), whereas the majority of cells migrated away from the other cell (outcomes 2 and 3 in Fig. 3a-e of [22]) when $c=12.4$ nM (see right bars in Fig. 3e of [22]).
Appendix D Experimental methods
D.1 Cells and reagents
KIC is a murine pancreatic cancer cell line derived from a genetically engineered murine pancreatic adenocarcinoma, featuring a Kras mutation along with Ink4a locus (Ink4a/ArfL/L) deletion [43, 44, 45]. In response to transforming growth factor-$\beta$ (TGF-$\beta$) stimulation, KIC cells exhibit a mesenchymal phenotype, enhanced invasive capacity, and directed migration [14, 15, 46]. KIC was cultured in RPMI 1640 supplemented with 1% penicillin/streptomycin (P/S) and 5% fetal bovine serum (FBS). Cells were routinely harvested using 0.05% trypsin and 0.53 mM EDTA (Life Technologies, CA, USA) upon reaching approximately 80% confluency in 25 cm2 T-flasks. They were incubated at 37C with 5% CO2. The harvested cells were either used for experiments or sub-cultured, ensuring they remained below the 19th passage.
D.2 Microfluidic platform
The microfluidic device consists of two source channels for chemical gradients and one central channel for culturing cells with 100 $\mu$m thickness [44, 14, 15]. As shown in Fig. 4(a), source 1 and source 2 channels were connected to the reservoir to perfuse the medium into the central channel. The central channel width is 1 mm, which is filled with 2 mg/ml collagen matrix with cells. The concentration of TGF-$\beta$ (Human TGF-$\beta$1, Thermo Fisher) was controlled at source 1 and source 2 channels to generate chemical gradients in the central channel. To produce a 5 nM/mm TGF-$\beta$ gradient, the source 2 channel was filled with $c_{2}=5$ nM TGF-$\beta$, while the source 1 channel was filled with cell culture medium without TGF-$\beta$ ($c_{1}=0$). To create a small TGF-$\beta$ gradient with flow, the source 2 channel was filled with $c_{2}=5$ nM TGF-$\beta$, and the source 1 channel was filled with $c_{1}=2.5$ nM TGF-$\beta$. To create a large TGF-$\beta$ gradient with flow, the source 2 channel was filled with $c_{2}=10$ nM TGF-$\beta$, and the source 1 channel was filled with cell culture medium without TGF-$\beta$ ($c_{1}=0$). The initial cell density was maintained at $7\times 10^{5}$ cells per ml across all groups. After loading the cells into the microfluidic chip, the cells were cultured in the regular medium for 24 hours. Subsequently, they were exposed to the engineered signaling environment under experimental conditions. To control the flow velocity at approximately 2.5 $\mu$m/s, we applied a pressure difference between two source channels [15].
D.3 Characterization of cell migration
Live-cell time-lapse imaging was performed using an inverted microscope with Zeiss incubation system (Zeiss AXIO Observer 7, Germany) at 37C with 5% CO2. KIC cells were imaged every 15 minutes for 3 hours and cell trajectories were analyzed using ImageJ. In gathering cell trajectories, we excluded trajectories of dividing cells and stationary cells. Dividing cells were excluded to prevent effects on cell polarity, while stationary cells were excluded to avoid underestimating cell movement characteristics. Directed cell migration is defined by motility and directional accuracy. The direction of cell trajectories was evaluated relative to environmental signals. Directional accuracy was quantified using the directional accuracy index, DAI $=\cos\theta$, where $\theta$ represents the angle between a trajectory’s net displacement and the direction of the environmental cue [Fig. 4(b)]. The displacement is defined by a straight line connecting the trajectory’s initial and final points. The DAI ranges from $-1$ to $1$. A DAI of 1 signifies that the cell moves precisely in the direction of the environmental signal, while a DAI of 0 indicates random motion. Conversely, a DAI of $-1$ means the cell moves in the exact opposite direction of the signal. A more detailed explanation of DAI can be found in previous studies [44, 15].
D.4 Analysis of DAI distributions
The distributions in Fig. 4(g) were analyzed using the one-sample Wilcoxon ($p_{\rm W}$) test and the nonparametric Mann-Whitney test ($p_{\rm MW}$) for statistical comparison. Both the TGF-$\beta$-only condition ($p_{\rm W}<0.0001$, $p_{\rm MW}<0.0001$) and the flow-only condition ($p_{\rm W}<0.01$, $p_{\rm MW}=0.0862$) are significantly different from the control. As in [15], the small- and large-gradient conditions with flow are significantly different from each other ($p_{\rm W}=0.0247$, $p_{\rm MW}=0.0145$).